# One Tool Is Enough: Reinforcement Learning for Repository-Level LLM Agents

### v3

@@ 1 Introduction
With the rapid advancement of Large Language Models (LLMs) (liu2024deepseek; team2024qwen2; yang2025qwen3), equipping LLMs with pre-built tools to form LLM agents has become a common paradigm for expanding their capabilities (shen2024llm; yuan2024easytool; lu2024toolsandbox). In the domain of software engineering (SWE), although LLM agents can effectively handle simple programming tasks (hui2024qwen2; guo2024deepseek), their ability to operate on large-scale open-source software (OSS) repositories remains limited. SWE-bench (jimenez2023swe) currently serves as the most comprehensive benchmark for evaluating whether LLMs can resolve real-world GitHub issues. All pretrained LLMs can not process the whole repository directly due to context limits. While SWE-Agent (jimenez2023swe) provides moderate gains, it remains far from enabling robust repository-level reasoning.
@@
Most existing agents rely on test-time scaling applied directly to pretrained LLMs (liu2023dynamic; chen-etal-2025-locagent; schmidgall2025agent). In software engineering (SWE) tasks, tools are essential rather than optional: real-world repositories are far larger than the context window of current LLMs, making it impossible to process an entire codebase in a single forward pass. Agents must therefore iteratively invoke tools to retrieve partial information from the repository and interleave natural-language reasoning with tool calls.
@@
However, mainstream LLM agents (chen-etal-2025-locagent; liu2025codexgraph; xiang2025promptsculptor; wang2023promptagent; chen2024comm) are rarely exposed to such agentic interaction patterns during pretraining and typically acquire tool usage only through few-shot prompting. Such in-context demonstrations are insufficient for learning complex multi-step tool-chaining behaviors, especially under limited context windows. Moreover, because tool definition spaces are effectively unbounded, pretrained models cannot fully internalize their semantics without post-training. To mitigate these issues, post-training paradigms such as Supervised Finetuning (SFT) (ma2025toolintegratedreinforcementlearningrepo) and Reinforcement Learning with Verifiable Rewards (RLVR) (yu2025dapo; yue2025vapo) have been applied, with promising results in domains including retrieval agents (jin2025search), GUI agents (hong2024cogagent), and math agents (yan2025mathagent).
@@
Directly training an agent to fix software issues, however, remains difficult. A single bug often admits multiple valid patches, making string-level evaluation unreliable. The only precise evaluation method requires executing candidate patches inside a dedicated Docker environment for each repository (deepswe2025), which is prohibitively expensive. To make training more tractable, we adopt a simplified yet widely generalizable assignment: issue localization. Prior work shows that a software issue becomes substantially easier to resolve once the relevant functions and files are correctly identified (chen-etal-2025-locagent; ma2025toolintegratedreinforcementlearningrepo; xia2024agentless; jiang2025cosil). Since modern OSS repositories contain a significant amount of code—far beyond any LLM’s context window—localization drastically reduces the search space and improves downstream solvability. Crucially, localization outputs a discrete set of paths, enabling verifiable, string-level evaluation that is compatible with scalable training frameworks such as SFT and RLVR.
@@
Existing localization agents (ma2025toolintegratedreinforcementlearningrepo; chen-etal-2025-locagent; SWESwiss2025) typically rely on multiple tools, including SearchClass, SearchMethods, and GetImports. Although effective to some extent, these tools considers high-level abstractions (classes, function, etc) of programing languages, which do not reflect how code actually executes. High-level abstractions, such as classes or inheritance, disappear after compilation, leaving only sequential execution and jump operations. Since modern LLMs already excel at modeling sequential dependencies, we focus on enhancing their ability to jump across the repository—that is, to follow and inspect the source definition of symbols as they appear in execution. To this end, we introduce a single, structurally grounded tool: jump, which retrieves the precise definition of a given symbol. Details of this tool are provided in Sec. 3.3.
@@ 2.1 Agentic Training
LLM agents are promising methods to equip models with complex tools while reasoning (li2024personal; huang2024understanding; guo2024large). However, because most pretrained LLMs are trained on texts only and developers can define any tools, most tools are out-of-domain (OOD) for LLMs. Even for the most powerful models, failures often happen when calling the new-defined tools due to wrong calling format or failed parameter parsing. Thus, training a LLM to master new-defined tool is critical for LLM agents. Intuitively, the tool-calling trajectories can be generated by a more powerful LLM, and such trajectories can be used to train a student model via supervised finetuning (SFT) (chen-etal-2025-locagent). However, this pipeline requires a stronger teacher model which has capability to master the tool. Recently, more methods have emerged with no teacher-model required. Rejected-sampled finetuning (RFT) (ahn2024large) utilizes generated trajectories of the agent itself via multiple rollouts. Agentic RL (jin2025search) is an on-policy RLVR methods requiring only the result for verifiying trajectories. Such training methods yield remarkable results when the tools are search engines (jin2025search), python executer (jimenez2023swe), calculator (yan2025mathagent), and visual models (gupta2023visual).
@@ 2.2 Software Engineering Agents
The introduction of SWE-bench (jimenez2023swe; yang2024swe) has motivated a range of agentic pipelines for software engineering (SWE) tasks. Among them, SWE-agent (yang2024sweagent) and OpenHands (wang2025openhands) are widely adopted frameworks that equip agents with tools for interacting with computing environments (terminal in Linux). Workflow-based methods such as Agentless (xia2024agentless) decomposes issue resolution into localization, repair, and validation subproblems. PatchPilot (li2025patchpilot) adds reproduction and refinement process in the workflow. CORTEXA (sohrabizadehnemotron) reflects the repository into an embedding space to enable efficient retrieval. chen-etal-2025-locagent builds the respository as a graph and applied graph-level searching tools for localization, and wang2025openhands furthermore integrated commit history as agent memory. RepoLens (wang2025extracting) equip conceptual information of the respository to enable repo-level understanding. These pipelines are training-free, and yield competitive results.
To enable task-specific training, DeepSWE (deepswe2025) and SWE-Swiss (SWESwiss2025) employ reinforcement learning and achieve strong performance. SWE-Gym (pan2024training) builds a virtual environment to train an agent for this task. However, end-to-end training remains costly because patch evaluation requires executing Docker containers across numerous repositories, and the Docker container hinders guided search methods due to its non-serializablity (zainullina2025guided). Consequently, issue localization has emerged as a computationally efficient alternative, aiming to identify faulty components—at file or function level—rather than generating full patches.
Recent localization agents include LocAgent (chen-etal-2025-locagent) and CoSIL (jiang2025cosil), which model codebases as graphs and integrates them into LLMs, and OrcaLoca (yu2025orcalocallmagentframework), which enhances efficiency through priority scheduling, action decomposition, and context pruning. From an open-source perspective, RepoSearcher (ma2025toolintegratedreinforcementlearningrepo), trained with distillation and RL on the Qwen model family (team2024qwen2), represents a notable advancement.
Nevertheless, prior agents overlook the structural relations within repositories—where modules, classes, and functions are cross-referenced across files—and typically rely on multiple search tools for symbol definition retrieval, amplifying error propagation (see Sec. 3). In contrast, we employ a single execution-logic-focused tool, reducing usage complexity. Finally, our approach constitutes the first localization agent trained directly from pretrained models, without relying on distillation-based supervised finetuning, a crucial stage in both RepoSearcher (ma2025toolintegratedreinforcementlearningrepo) and LocAgent (chen-etal-2025-locagent).
@@ 3.5 Reinforcement Learning
We apply reinforcement learning with verifiable rewards to train the agent directly from the pretrained model. We do not require a more powerful teacher model for distillation as a warmup, which is the technique of (ma2025toolintegratedreinforcementlearningrepo). In practice, we apply Group Reference Policy Optimization (GRPO), which has the loss function:
@@ 4.1 Experimnent Setup
We extract valid samples from SWE-smith (yang2025swe) to form the training set. We apply Qwen2.5-7B-Instruct with RepoNavigator to sample each data for 16 times. A sample is abandoned if all 16 scores are zero. For validation, we test our method on SWE-bench-verified (jimenez2023swe), which is a human-verified subset of SWE-bench. We additionally test our method on a subset of SWE-bench-pro (yang2025swe) (which is a new and more difficult benchmark) for generalization. For ground-truth locations, we directly use the locations in golden patches. All datasets are open-source and are built on real-world github issues.
@@
Previous works (chen-etal-2025-locagent; ma2025toolintegratedreinforcementlearningrepo) applied recall and precision as metrics. However, because the predicted locations and ground-truth locations are sets of strings, recall and precision singularly can not reflect the performance fairly. Thus, we utilize sample-F1 (which is the averaged score of per-sample F1 values) and IoU (intersection out of union) as our core metrics. At the same time, we also present the recall and precision scores to align with previous methods, although they do not reflect the methods’ performance fairly.
@@
For the 7B model, we conduct GRPO with 8 Tesla-A100-80G GPUs. For the 14B and 32B model, we train it with 16 Tesla-A100-80G GPUs. We apply verl (shen2024llm) as the training framework, and we apply vLLM (kwon2023efficient) as the inference engine. We train the model for 1 epoch, while the training batch size is fixed to 128 on 4k training samples filtered from SWE-smith, with maximum prompt length and max response length both set to 10240. Additionally, we rollout 8 times for each sample, and the temperature is set to 1.0 to encourage exploration. We use greedy decoding in the inference stage to ensure stable performance. More implementation details are provided in Appendix. B.
@@ 4.2 Effectiveness
We compare our method against Locagent (chen-etal-2025-locagent), CoSIL (jiang2025cosil), Agentless (xia2024agentless), Orcaloca (yu2025orcalocallmagentframework), and RepoSearcher (ma2025toolintegratedreinforcementlearningrepo). Detailed explaination of baseline methods are presented in Appendix. A.
@@ Results
Compared with RepoSearcher, which is distilled from claude-3.7-sonnet (anthropic2025claude37sonnet) and reinforced by GRPO, trained RepoNavigator outperforms it on all metrices except recall. Moreover, we found that our training-free method outperforms RepoSearcher for 14B models. This is probably due to the simplified tool we integrate to the agent (see Sec. 5 for more details).
To assess the generalizability of RepoNavigator, we present its performance on Python samples from the SWE-bench-Pro dataset (yang2025swe) in Table 2. The results on this dataset are consistent with those observed on SWE-bench_Verified. While we cannot fully exclude the potential influence of data leakage in SWE-bench_Verified, we can make a stronger claim regarding SWE-bench_Pro, as it was released after the publication of the Qwen2.5 series.
@@ 4.3 Training Strategy Comparison
To explore the capability of GRPO on agentic training, we compare GRPO against RFT-only and RFT+GRPO. As presented in Fig. 3, directly training with GRPO outperformes RFT-only and RFT+GRPO. Moreover, although RFT has accetable performance, the more steps RFT proceeds, the less improvement GRPO makes after the cold start. This conclusion contradicts with previous SWE agents trained with RL (ma2025toolintegratedreinforcementlearningrepo), however, it aligns with the broader field of reinforcement learning, where RFT and SFT (as a cold start) is effective only when the pretrained model is not strong enough (guo2024deepseek). When the pretrained model is strong enough and data is high-quality, directly training a model with RL is better than training after SFT (RFT) as its cold start.
@@ 5.4 Verification
To further verify this proposal, we change the tool set of RepoNavigator and conduct RL training with only the outcome reward. We add excessive tools which were frequently used in previous works (chen-etal-2025-locagent; ma2025toolintegratedreinforcementlearningrepo; jiang2025cosil) and present the result in Table. 4. GetClass/GetFunc takes a class/function name as input and outputs the class/function definition. GetStruc takes no input and outputs the repository’s structure. The results clearly implies that additional tools do not increase model’s performance. This inspires researchers to develop less but more capable tools.
@@ 6 Conclusion
Table 5: Illustration of the tool sets of agentic methods. Among all methods, only RepoNavigator takes the jump logic into consideration, and RepoNavigator has the smallest number of tools.
Method Number Tools Input Output
CoSIL 3 search_class_node file_path, class_name The code snippet of a class node.
search_class_function_node file_path, class_name, function_name Get the code snippet of a class member function node.
search_file_function_node file_path, function_name Get the code snippet of a static function node.
Locagent 3 SearchEntity Keywords Related entities with code snippets.
TraverseGraph Start entity ids, direction, traverse hops, entity types, relation types A traversed subgraph with entities and relations.
RetrieveEntiry Entity ids The complete code of a specific entity.
Orcaloca 5 search_file_contents file_name, directory_path The file contents or skeleton (if exceeds 200 lines).
search_class class_name, method_name The class contents or skeleton (if exceeds 200 lines).
search_method_in_class class_name, method_name, file_path The method code snippet.
search_callable query_name, file_path The matched code snippet.
search_source_code file_path, source_code The related function/class code snippet.
RepoSearcher 5 GetRepoStructure None The repository file structure.
GetImportOfFile file, class The imports of file.
SearchClass file, class Code content of the searched class.
SearchFunction file, function Code content of the searched function.
SearchClassMethod file, class, method Code content of the searched method.
Ours 1 Jump file_path, symbol, index (optional) Definition code snippet of the symbol.
@@ Appendix A Detailed Illustration of Baselines
Agentless (xia2024agentless) is a workflow for issue localization. First, it identifies suspicious files in the repository. Second, relevant classes and functions are detected. Third, precise locations for edit are given by LLMs based on the classes and functions.
@@
CoSIL (jiang2025cosil) is an agent which first conduct file-level localization and then conduct function-level localization. CoSIL dynamically constructs call graphs of modules (class, functions) during the repo-level searching process, and applies context pruning to effectively reduce the searching scope.
@@
LocAgent (chen-etal-2025-locagent) is almost a fully-automatic LLM agent besides its planning prompt concatenated into the context at the beginning of the searching process. It builds the whole repository into a direct heterogeneous graph, whose nodes are files, classes, and functions. Additionally, edges are built by dependencies such as imports and invocations. Multiple graph-level searching tools are equipped to the LLM for multi-hop reasoning.
@@
OrcaLoca (yu2025orcalocallmagentframework) is an LLM-based agent framework for software issue localization that performs repository-level exploration via priority-based action scheduling. It incrementally decomposes file- and class-level searches into fine-grained function-level actions with relevance scoring, while applying distance-aware context pruning on a code graph to efficiently narrow down potential bug locations. Because we failed to reproduce the code of Orcaloca, we did not evaluate it on SWE-ben_Pro, and we directly report the results (provided by jiang2025cosil) on SWE-bench_Verified.
@@
RepoSearcher (ma2025toolintegratedreinforcementlearningrepo) is an agent that first conducts file-level localization and then function-level localization, which aligns with CoSIL. RepoSearcher introduced the first training framework ToolTrain for localization agents, which is composed of distilling from a close-source model (Claude3.7-Sonnet in RepoSeacher) as warmup and reinforcement learning to further enhance the performance.

### v4

@@ 1 Introduction
With the rapid advancement of Large Language Models (LLMs) (Liu et al., 2024; Team, 2024; Yang et al., 2025a), equipping LLMs with pre-built tools to form LLM agents has become a common paradigm for expanding their capabilities (Shen, 2024; Yuan et al., 2024; Lu et al., 2024). In the domain of software engineering (SWE), although LLM agents can effectively handle simple programming tasks (Hui et al., 2024; Guo et al., 2024a), their ability to operate on large-scale open-source software (OSS) repositories remains limited. SWE-bench (Jimenez et al., 2023) currently serves as the most comprehensive benchmark for evaluating whether LLMs can resolve real-world GitHub issues. All pretrained LLMs can not process the whole repository directly due to context limits. While SWE-Agent (Jimenez et al., 2023) provides moderate gains, it remains far from enabling robust repository-level reasoning.
@@
Most existing agents rely on test-time scaling applied directly to pretrained LLMs (Liu et al., 2023; Chen et al., 2025; Schmidgall et al., 2025). In software engineering (SWE) tasks, tools are essential rather than optional: real-world repositories are far larger than the context window of current LLMs, making it impossible to process an entire codebase in a single forward pass. Agents must therefore iteratively invoke tools to retrieve partial information from the repository and interleave natural-language reasoning with tool calls.
@@
However, mainstream LLM agents (Chen et al., 2025; Liu et al., 2025; Xiang et al., 2025; Wang et al., 2023; Chen et al., 2024) are rarely exposed to such agentic interaction patterns during pretraining and typically acquire tool usage only through few-shot prompting. Such in-context demonstrations are insufficient for learning complex multi-step tool-chaining behaviors, especially under limited context windows. Moreover, because tool definition spaces are effectively unbounded, pretrained models cannot fully internalize their semantics without post-training. To mitigate these issues, post-training paradigms such as Supervised Finetuning (SFT) (Ma et al., 2025) and Reinforcement Learning with Verifiable Rewards (RLVR) (Yu et al., 2025a; Yue et al., 2025) have been applied, with promising results in domains including retrieval agents (Jin et al., 2025), GUI agents (Hong et al., 2024), and math agents (Yan et al., 2025).
@@
Directly training an agent to fix software issues, however, remains difficult. A single bug often admits multiple valid patches, making string-level evaluation unreliable. The only precise evaluation method requires executing candidate patches inside a dedicated Docker environment for each repository (Luo et al., 2025), which is prohibitively expensive. To make training more tractable, we adopt a simplified yet widely generalizable assignment: issue localization. Prior work shows that a software issue becomes substantially easier to resolve once the relevant functions and files are correctly identified (Chen et al., 2025; Ma et al., 2025; Xia et al., 2024; Jiang et al., 2025). Since modern OSS repositories contain a significant amount of code—far beyond any LLM’s context window—localization drastically reduces the search space and improves downstream solvability. Crucially, localization outputs a discrete set of paths, enabling verifiable, string-level evaluation that is compatible with scalable training frameworks such as SFT and RLVR.
@@
Existing localization agents (Ma et al., 2025; Chen et al., 2025; He et al., 2025) typically rely on multiple tools, including SearchClass, SearchMethods, and GetImports. Although effective to some extent, these tools considers high-level abstractions (classes, function, etc) of programing languages, which do not reflect how code actually executes. High-level abstractions, such as classes or inheritance, disappear after compilation, leaving only sequential execution and jump operations. Since modern LLMs already excel at modeling sequential dependencies, we focus on enhancing their ability to jump across the repository—that is, to follow and inspect the source definition of symbols as they appear in execution. To this end, we introduce a single, structurally grounded tool: jump, which retrieves the precise definition of a given symbol. Details of this tool are provided in Sec. 3.3.
@@ 2.1 Agentic Training
LLM agents are promising methods to equip models with complex tools while reasoning (Li et al., 2024; Huang et al., 2024; Guo et al., 2024b). However, because most pretrained LLMs are trained on texts only and developers can define any tools, most tools are out-of-domain (OOD) for LLMs. Even for the most powerful models, failures often happen when calling the new-defined tools due to wrong calling format or failed parameter parsing. Thus, training a LLM to master new-defined tool is critical for LLM agents. Intuitively, the tool-calling trajectories can be generated by a more powerful LLM, and such trajectories can be used to train a student model via supervised finetuning (SFT) (Chen et al., 2025). However, this pipeline requires a stronger teacher model which has capability to master the tool. Recently, more methods have emerged with no teacher-model required. Rejected-sampled finetuning (RFT) (Ahn et al., 2024) utilizes generated trajectories of the agent itself via multiple rollouts. Agentic RL (Jin et al., 2025) is an on-policy RLVR methods requiring only the result for verifiying trajectories. Such training methods yield remarkable results when the tools are search engines (Jin et al., 2025), python executer (Jimenez et al., 2023), calculator (Yan et al., 2025), and visual models (Gupta and Kembhavi, 2023).
@@ 2.2 Software Engineering Agents
The introduction of SWE-bench (Jimenez et al., 2023; Yang et al., 2024b) has motivated a range of agentic pipelines for software engineering (SWE) tasks. Among them, SWE-agent (Yang et al., 2024a) and OpenHands (Wang et al., 2025a) are widely adopted frameworks that equip agents with tools for interacting with computing environments (terminal in Linux). Workflow-based methods such as Agentless (Xia et al., 2024) decomposes issue resolution into localization, repair, and validation subproblems. PatchPilot (Li et al., 2025) adds reproduction and refinement process in the workflow. CORTEXA (Sohrabizadeh et al., ) reflects the repository into an embedding space to enable efficient retrieval. Chen et al. (2025) builds the respository as a graph and applied graph-level searching tools for localization, and Wang et al. (2025a) furthermore integrated commit history as agent memory. RepoLens (Wang et al., 2025b) equip conceptual information of the respository to enable repo-level understanding. These pipelines are training-free, and yield competitive results.
To enable task-specific training, DeepSWE (Luo et al., 2025) and SWE-Swiss (He et al., 2025) employ reinforcement learning and achieve strong performance. SWE-Gym (Pan et al., 2024) builds a virtual environment to train an agent for this task. However, end-to-end training remains costly because patch evaluation requires executing Docker containers across numerous repositories, and the Docker container hinders guided search methods due to its non-serializablity (Zainullina et al., 2025). Consequently, issue localization has emerged as a computationally efficient alternative, aiming to identify faulty components—at file or function level—rather than generating full patches.
Recent localization agents include LocAgent (Chen et al., 2025) and CoSIL (Jiang et al., 2025), which model codebases as graphs and integrates them into LLMs, and OrcaLoca (Yu et al., 2025b), which enhances efficiency through priority scheduling, action decomposition, and context pruning. From an open-source perspective, RepoSearcher (Ma et al., 2025), trained with distillation and RL on the Qwen model family (Team, 2024), represents a notable advancement.
Nevertheless, prior agents overlook the structural relations within repositories—where modules, classes, and functions are cross-referenced across files—and typically rely on multiple search tools for symbol definition retrieval, amplifying error propagation (see Sec. 3). In contrast, we employ a single execution-logic-focused tool, reducing usage complexity. Finally, our approach constitutes the first localization agent trained directly from pretrained models, without relying on distillation-based supervised finetuning, a crucial stage in both RepoSearcher (Ma et al., 2025) and LocAgent (Chen et al., 2025).
@@ 3.5 Reinforcement Learning
We apply reinforcement learning with verifiable rewards to train the agent directly from the pretrained model. We do not require a more powerful teacher model for distillation as a warmup, which is the technique of (Ma et al., 2025). In practice, we apply Group Reference Policy Optimization (GRPO), which has the loss function:
@@ 4.1 Experimnent Setup
We extract valid samples from SWE-smith (Yang et al., 2025b) to form the training set. We apply Qwen2.5-7B-Instruct with RepoNavigator to sample each data for 16 times. A sample is abandoned if all 16 scores are zero. For validation, we test our method on SWE-bench-verified (Jimenez et al., 2023), which is a human-verified subset of SWE-bench. We additionally test our method on a subset of SWE-bench-pro (Yang et al., 2025b) (which is a new and more difficult benchmark) for generalization. For ground-truth locations, we directly use the locations in golden patches. All datasets are open-source and are built on real-world github issues.
@@
Previous works (Chen et al., 2025; Ma et al., 2025) applied recall and precision as metrics. However, because the predicted locations and ground-truth locations are sets of strings, recall and precision singularly can not reflect the performance fairly. Thus, we utilize sample-F1 (which is the averaged score of per-sample F1 values) and IoU (intersection out of union) as our core metrics. At the same time, we also present the recall and precision scores to align with previous methods, although they do not reflect the methods’ performance fairly.
@@
For the 7B model, we conduct GRPO with 8 Tesla-A100-80G GPUs. For the 14B and 32B model, we train it with 16 Tesla-A100-80G GPUs. We apply verl (Shen, 2024) as the training framework, and we apply vLLM (Kwon et al., 2023) as the inference engine. We train the model for 1 epoch, while the training batch size is fixed to 128 on 4k training samples filtered from SWE-smith, with maximum prompt length and max response length both set to 10240. Additionally, we rollout 8 times for each sample, and the temperature is set to 1.0 to encourage exploration. We use greedy decoding in the inference stage to ensure stable performance. More implementation details are provided in Appendix. B.
@@ 4.2 Effectiveness
We compare our method against Locagent (Chen et al., 2025), CoSIL (Jiang et al., 2025), Agentless (Xia et al., 2024), Orcaloca (Yu et al., 2025b), and RepoSearcher (Ma et al., 2025). Detailed explaination of baseline methods are presented in Appendix. A.
@@ Results
Compared with RepoSearcher, which is distilled from claude-3.7-sonnet (Anthropic, 2025) and reinforced by GRPO, trained RepoNavigator outperforms it on all metrices except recall. Moreover, we found that our training-free method outperforms RepoSearcher for 14B models. This is probably due to the simplified tool we integrate to the agent (see Sec. 5 for more details).
To assess the generalizability of RepoNavigator, we present its performance on Python samples from the SWE-bench-Pro dataset (Yang et al., 2025b) in Table 2. The results on this dataset are consistent with those observed on SWE-bench_Verified. While we cannot fully exclude the potential influence of data leakage in SWE-bench_Verified, we can make a stronger claim regarding SWE-bench_Pro, as it was released after the publication of the Qwen2.5 series.
@@ 4.3 Training Strategy Comparison
To explore the capability of GRPO on agentic training, we compare GRPO against RFT-only and RFT+GRPO. As presented in Fig. 3, directly training with GRPO outperformes RFT-only and RFT+GRPO. Moreover, although RFT has accetable performance, the more steps RFT proceeds, the less improvement GRPO makes after the cold start. This conclusion contradicts with previous SWE agents trained with RL (Ma et al., 2025), however, it aligns with the broader field of reinforcement learning, where RFT and SFT (as a cold start) is effective only when the pretrained model is not strong enough (Guo et al., 2024a). When the pretrained model is strong enough and data is high-quality, directly training a model with RL is better than training after SFT (RFT) as its cold start.
@@ 5.4 Verification
To further verify this proposal, we change the tool set of RepoNavigator and conduct RL training with only the outcome reward. We add excessive tools which were frequently used in previous works (Chen et al., 2025; Ma et al., 2025; Jiang et al., 2025) and present the result in Table. 4. GetClass/GetFunc takes a class/function name as input and outputs the class/function definition. GetStruc takes no input and outputs the repository’s structure. The results clearly implies that additional tools do not increase model’s performance. This inspires researchers to develop less but more capable tools.
@@ Appendix A Detailed Illustration of Baselines
Agentless (Xia et al., 2024) is a workflow for issue localization. First, it identifies suspicious files in the repository. Second, relevant classes and functions are detected. Third, precise locations for edit are given by LLMs based on the classes and functions.
@@
CoSIL (Jiang et al., 2025) is an agent which first conduct file-level localization and then conduct function-level localization. CoSIL dynamically constructs call graphs of modules (class, functions) during the repo-level searching process, and applies context pruning to effectively reduce the searching scope.
@@
LocAgent (Chen et al., 2025) is almost a fully-automatic LLM agent besides its planning prompt concatenated into the context at the beginning of the searching process. It builds the whole repository into a direct heterogeneous graph, whose nodes are files, classes, and functions. Additionally, edges are built by dependencies such as imports and invocations. Multiple graph-level searching tools are equipped to the LLM for multi-hop reasoning.
@@
OrcaLoca (Yu et al., 2025b) is an LLM-based agent framework for software issue localization that performs repository-level exploration via priority-based action scheduling. It incrementally decomposes file- and class-level searches into fine-grained function-level actions with relevance scoring, while applying distance-aware context pruning on a code graph to efficiently narrow down potential bug locations. Because we failed to reproduce the code of Orcaloca, we did not evaluate it on SWE-ben_Pro, and we directly report the results (provided by Jiang et al. (2025)) on SWE-bench_Verified.
@@
RepoSearcher (Ma et al., 2025) is an agent that first conducts file-level localization and then function-level localization, which aligns with CoSIL. RepoSearcher introduced the first training framework ToolTrain for localization agents, which is composed of distilling from a close-source model (Claude3.7-Sonnet in RepoSeacher) as warmup and reinforcement learning to further enhance the performance.